# Closure Discovery for Coarse-Grained Partial Differential Equations using Multi-Agent Reinforcement Learning

### v1

@@ Abstract
Reliable predictions of critical phenomena, such as weather, wildfires and epidemics are often founded on models described by Partial Differential Equations (PDEs). However, simulations that capture the full range of spatio-temporal scales in such PDEs are often prohibitively expensive. Consequently, coarse-grained simulations that employ heuristics and empirical closure terms are frequently utilized as an alternative. We propose a novel and systematic approach for identifying closures in under-resolved PDEs using Multi-Agent Reinforcement Learning (MARL). The MARL formulation incorporates inductive bias and exploits locality by deploying a central policy represented efficiently by Convolutional Neural Networks (CNN). We demonstrate the capabilities and limitations of MARL through numerical solutions of the advection equation and the Burgers’ equation. Our results show accurate predictions for in- and out-of-distribution test cases as well as a significant speedup compared to resolving all scales.
Multi-Agent Reinforcement Learning, Closure Modeling, Partial Differential Equations, Coarse-Graining
@@ 1 Introduction
Simulations of critical phenomena such as climate, ocean dynamics and epidemics, have become essential for decision-making, and their veracity, reliability, and energy demands have great impact on our society. Many of these simulations are based on models described by PDEs expressing system dynamics that span multiple spatio-temporal scales. Examples include turbulence (Wilcox, 1988), neuroscience (Dura-Bernal et al., 2019), climate (Council, 2012) and ocean dynamics (Mahadevan, 2016). Today, we benefit from decades of remarkable efforts in the development of numerical methods, algorithms, software, and hardware and witness simulation frontiers that were unimaginable even a few years ago (Hey, 2009). Large-scale simulations that predict the system’s dynamics may use trillions of computational elements(Rossinelli et al., 2013) to resolve all spatio-temporal scales, but these often address only idealized systems and their computational cost prevents experimentation and uncertainty quantification. On the other hand, reduced order and coarse-grained models are fast, but limited by the linearization of complex system dynamics while their associated closures are often based on heuristics (Peng et al., 2021).
We propose to complement coarse-grained simulations with closures in the form of a policies that are discovered by a multi-agent reinforcement learning framework. A central policy that is based on CNNs and uses locality as inductive bias, corrects the error of the numerical discretization and improves the overall accuracy. This approach is inspired by recent work in Reinforcement Learning (RL) for image reconstruction (Furuta et al., 2019). The agents learn and act locally on the grid, in a manner that is reminiscent of the numerical discretizations of PDEs based on Taylor series approximations. Hence, each agent only gets information from its spatial neighborhood. Together with a central policy, this results in efficient training, as each agent needs to process only a limited amount of observations. After training, the framework is capable of accurate predictions in both in- and out-of-distribution test cases. We remark that the actions taken by the agents are highly correlated with the numerical errors and can be viewed as an implicit correction to the effective convolution performed via the coarse-grained discretization of the PDEs (Bergdorf et al., 2005).
@@ 2 Related Work
The main contribution of this work is a MARL algorithm for the discovery of closures for coarse-grained PDEs. The algorithm provides an automated process for the correction of the errors of the associated numerical discretization. We show that the proposed method is able to capture scales beyond the training regime and provides a potent method for solving PDEs with high accuracy and limited resolution.
@@
Machine Learning and Partial Differential Equations: In recent years, there has been significant interest in learning the solution of PDEs using Neural Networks. Techniques such as PINNs (Raissi et al., 2019; Karniadakis et al., 2021), DeepONet (Lu et al., 2021), the Fourier Neural Operator (Li et al., 2020), NOMAD (Seidman et al., 2022), Clifford Neural Layers (Brandstetter et al., 2023) and an invertible formulation (Kaltenbach et al., 2023) have shown promising results for both forward and inverse problems. However, there are concerns about their accuracy and related computational cost, especially for low-dimensional problems (Karnakov et al., 2022). These methods aim to substitute numerical discretizations with neural nets, in contrast to our RL framework, which aims to complement them. Moreover, their loss function is required to be differentiable, which is not necessary for the stochastic formulation of the RL reward.
Reinforcement Learning: Our work is based upon recent advances in RL (Sutton & Barto, 2018) including PPO (Schulman et al., 2017) and the Remember and Forget Experience Replay (ReFER) algorithm (Novati et al., 2019). It is related to other methods that employ multi-agent learning (Novati et al., 2021; Albrecht et al., 2023; Wen et al., 2022; Bae & Koumoutsakos, 2022; Yang & Wang, 2020) for closures except that here we use as reward not global data (such as energy spectrum) but the local dicsretization error as quantified by fine-grained simulations. The present approach is designed to solve various PDEs using a central policy and it is related to similar work for image reconstruction (Furuta et al., 2019). We train a central policy for all agents in contrast to other methods where agents are trained on decoupled subproblems (Freed et al., 2021; Albrecht et al., 2023; Sutton et al., 2023).
@@
Closure Modeling: The development of machine learning methods for discovering closures for under-resolved PDEs has gained attention in recent years. Current approaches are mostly tailored to specific applications such as turbulence modeling (Ling et al., 2016; Durbin, 2018; Novati et al., 2021; Bae & Koumoutsakos, 2022) and use data such as energy spectra and drag coefficients of the flow in order to train the RL policies. In (Lippe et al., 2023), a more general framework based on diffusion models is used to improve the solutions of Neural Operators for temporal problems using a multi-step refinement process. Their training is based on supervised learning in contrast to the present RL approach which additionally complements existing numerical methods instead of neural network based surrogates (Li et al., 2020; Gupta & Brandstetter, 2022).
@@
Inductive Bias: The incorporation of prior knowledge regarding the physical processes described by the PDEs, into machine learning algorithms is critical for their training in the Small Data regime and for increasing the accuracy during extrapolative predictions (Goyal & Bengio, 2022). One way to achieve this is by shaping appropriately the loss function (Karniadakis et al., 2021; Kaltenbach & Koutsourelakis, 2020; Yin et al., 2021; Wang et al., 2021, 2022), or by incorporating parameterized mappings that are based on known constraints (Greydanus et al., 2019; Kaltenbach & Koutsourelakis, 2021; Cranmer et al., 2020). Our RL framework incorporates locality and is thus consistent with numerical discretizations that rely on local Taylor series based approximations. The incorporation of inductive bias in RL has also been attempted by focusing on the beginning of the training phase (Uchendu et al., 2023; Walke et al., 2023) in order to shorten the exploration phase.
@@ 3 Methodology
We consider a time-dependent, parametric, non-linear PDE defined on a regular domain $\Omega$. The solution of the PDE depends on its initial conditions (ICs) as well as its input parameters $C\in\mathcal{C}$. The PDE is discretized on a spatiotemporal grid and the resulting discrete set of equations is solved using numerical methods.
In turn, the number of the deployed computational elements and the structure of the PDE determine whether all of its scales have been resolved or whether the discretization amounts to a coarse-grained representation of the PDE. In the first case, the (fine grid simulation (FGS)) provides the discretized solution $\boldsymbol{\psi}$, whereas in (coarse grid simulation (CGS)) the resulting approximation is denoted by $\tilde{\boldsymbol{\psi}}$.111 In the following, all variables with the $\;\tilde{}\;$ are referring to the coarse-grid description. The RL policy can improve the accuracy of the solution $\tilde{\boldsymbol{\psi}}$ by introducing an appropriate forcing term in the right-hand side of the CGS. For this purpose, FGS of the PDE are used as training episodes and serve as the ground truth to facilitate a reward signal. The CGS and FGS employed herein are introduced in the next section. The proposed MARL framework is introduced in Section 3.2.
@@ 3.1 Coarse and Fine Grid Simulation
Similarly, we define the time stepping operator of the FGS as $\mathcal{F}:\Psi\times\mathcal{C}\rightarrow\Psi$. In the case of the CGS, $\tilde{C}\in\tilde{\mathcal{C}}$, is an adapted version of $C$, which for instance involves evaluating a parameter function using the coarse grid.
To simplify the notation, we set $n:={\tilde{n}}=d_{t}n_{f}$ in the following. Hence, we apply $\mathcal{F}$ $d_{T}$-times in order to describe the same time instant with the same index $n$ in both CGS and FGS. The update rule of the FGS are referred to as
@@
In line with the experiments performed in Section 4, and to further simplify the notation, we are dropping the third spatial dimension in the following presentation.
@@ 3.2 RL Environment
The environment of our RL framework is summarized in Figure 1. We define the state at step $n$ of the RL environment as the tuple $\boldsymbol{S}^{n}:=(\boldsymbol{\psi}^{n},\tilde{\boldsymbol{\psi}^{n}})$. This state is only partially observable as the policy is acting only in the CGS. The observation is defined as the coarse representation of the discretized solution function and the parameters $\boldsymbol{O}^{n}:=(\tilde{\boldsymbol{\psi}^{n}},\tilde{C}^{n})\in\mathcal{O}$. Our goal is to train a policy $\pi$ that makes the dynamics of the CGS to be close to the dynamics of the FGS. To achieve this goal, the action ${\boldsymbol{A}^{n}}\in\mathcal{A}:=\mathbb{R}^{k\times N_{x}\times N_{y}}$ at step $n$ of the environment is a collection of forcing terms for each discretization point of the CGS. In case the policy is later used to complement the CGS simulation the update function in Equation 1 changes to
@@
To encourage learning a policy that represents the non-resolved spatio-temporal scales, the reward is based on the difference between the CGS and FGS at time step $n$. In more detail, we define a local reward $\boldsymbol{R}^{n}\in\mathbb{R}^{\tilde{N}_{x}\times\tilde{N}_{y}}$ inspired by the reward proposed for image reconstruction in (Furuta et al., 2019):
@@
3.3 Multi-Agent Formulation
The policy $\boldsymbol{\pi}$ predicts a local action $\boldsymbol{A}^{n}_{i,j}\in\mathbb{R}^{k}$ at each discretization point which implies a very high dimensional continuous action space. Hence, formulating the closures with a single agentis very challenging. However, since the rewards are designed to be inherently local, locality can be used as inductive bias and the RL learning framework can be interpreted as a multi-agent problem (Furuta et al., 2019). One agent is placed at each discretization point of the coarse grid with a corresponding local reward $R_{ij}^{n}$. We remark that this approach augments adaptivity as one can place extra agents at additional, suitably selected, discretization points.
Each agent develops its own optimal policy and Equation 5 is replaced by
$\pi_{ij}^{*}=\underset{\pi_{ij}}{\operatorname{argmax}}\ \mathbb{E}_{\pi_{ij}}\left(\sum_{t=0}^{\infty}\gamma^{t}R_{ij}^{n}\right),$ (7)
Here, we used $\mathcal{O}(\tilde{N}_{x}\tilde{N}_{y})$ agents, which for typically used grid sizes used in numerical simulations, it becomes a large number compared to typical MARL problem settings (Albrecht et al., 2023).
We parametrize the local policies using neural networks. However, since training this many individual neural nets can become computationally expensive, we parametrize all the agents together using one fully convolutional network (FCN) (Long et al., 2015).
@@ 3.4 Parallelizing Local Policies with a FCN
We define the collection of policies in the FCN as
$\Pi^{FCN}:\mathcal{O}\rightarrow\mathcal{A}$ (8)
In further discussions, we will refer to $\Pi^{FCN}$ as the global policy. The policy $\pi_{ij}$ of the agent at point $(i,j)$ is subsequently implicitly defined through the global policy as
$\pi_{ij}(O_{ij}):=\left[\Pi^{FCN}(\boldsymbol{O})\right]_{:ij}.$ (9)
Here, $O_{ij}$ contains only a part of the input contained in $\boldsymbol{O}$  [^2] The exact content of $O_{ij}$ is depending on the receptive field of the FCN. For consistency, we will refer to $\boldsymbol{O}_{ij}$ as a local observation and $\boldsymbol{O}$ as the global observation. We note that the policies $\pi_{ij}(O_{ij})$ map the observation to a probability distribution at each discretization point (see Section 3.6 for details).
Similar to the global policy, the global value function is parametrized using a FCN as well. It maps the global observation to an expected return $H\in\mathcal{H}:=R^{\tilde{N}_{x}\times\tilde{N}_{y}}$ at each discretization point
$\mathcal{V}^{FCN}:\mathcal{O}\rightarrow\mathcal{H}.$ (10)
Similarly to the local policies, we define the local value function related to the agent at point $(i,j)$ as $v_{ij}(O_{ij}):=[\mathcal{V}^{FCN}(\boldsymbol{O})]_{ij}.$
3.5 Algorithmic Details
In order to solve the optimization problem in Equation 7 with the PPO algorithm (Schulman et al., 2017), we modify the single agent approach.
@@ 3.5 Algorithmic Details
$\mathbb{E}_{\boldsymbol{S},\boldsymbol{A}\sim\Pi^{FCN}}\left[\frac{1}{\tilde{N}_{x}\cdot\tilde{N}_{y}}\sum_{i,j=1}^{\tilde{N}_{x},\tilde{N}_{y}}L_{\pi_{ij}}(O_{ij},\boldsymbol{A}_{ij})\right]$ (11)
with the local version of the PPO objective $L_{\pi_{ij}}(O_{ij},\boldsymbol{A}_{ij})$. This corresponds to the local objective of the policy of the agent at point $(i,j)$ (Schulman et al., 2017).
@@
We have provided an overview of the modified PPO algorithm in Appendix A together with further details regarding the local version of the PPO objective $L_{\pi_{ij}}(O_{ij},\boldsymbol{A}_{ij})$. Our implementation of the adapted PPO algorithm is based on the single agent PPO algorithm of the Tianshou framework (Weng et al., 2022).
3.6 Neural Network Architecture
Figure 2: The IRCNN backbone takes the current global observation and maps it to a feature tensor. This feature tensor is passed into two different convolutional layers that predict the per-discretization point Gaussian parameters and state-values. In the case of a PDE with three spatial dimensions, the architecture would need to be based on three-dimensional convolutional layers instead.
The optimal policy is expected to compensate for errors introduced by the numerical method and its implementation on a coarse grid. We use the Image Restoration CNN (IRCNN) architecture proposed in (Zhang et al., 2017) as our backbone for the policy- and value-network. In Figure 2 we present an illustration of the architecture and show that the policy- and value network share the same backbone. The policy- and value network only differ by their last convolutional layer, which takes the features extracted by the backbone and maps them either to the parameters $\boldsymbol{\mu}_{A}\in\mathcal{A}$ and $\boldsymbol{\sigma}_{A}\in\mathcal{A}$ or the predicted return value for each agent. The local policies are assumed to be independent, such that we can write the distribution at a specific discretization point as
$\pi_{ij}(\boldsymbol{A}_{ij}|O_{ij})=\mathcal{N}(\boldsymbol{\mu}_{A,ij},\boldsymbol{\sigma}_{A,ij}).$ (12)
In the following, we are referring to the proposed MARL framework using a FCN for both the policy and the value function as CNN-based multi-agent reinforcement learning (CNN-MARL).

### v3

# Closure Discovery for Coarse-Grained Partial Differential Equations Using Grid-based Reinforcement Learning

@@ Abstract
Reliable predictions of critical phenomena, such as weather, wildfires and epidemics often rely on models described by Partial Differential Equations (PDEs). However, simulations that capture the full range of spatio-temporal scales described by such PDEs are often prohibitively expensive. Consequently, coarse-grained simulations are usually deployed that adopt various heuristics and empirical closure terms to account for the missing information. We propose a novel and systematic approach for identifying closures in under-resolved PDEs using grid-based Reinforcement Learning. This formulation incorporates inductive bias and exploits locality by deploying a central policy represented efficiently by a Fully Convolutional Network (FCN). We demonstrate the capabilities and limitations of our framework through numerical solutions of the advection equation and the Burgers’ equation. Our results show accurate predictions for in- and out-of-distribution test cases as well as a significant speedup compared to resolving all scales.
@@ 1 Introduction
Simulations of critical phenomena such as climate, ocean dynamics and epidemics, have become essential for decision-making, and their veracity, reliability, and energy demands have great impact on our society. Many of these simulations are based on models described by PDEs expressing system dynamics that span multiple spatio-temporal scales. Examples include turbulence [1], neuroscience [2], climate [3] and ocean dynamics [4].
Today, we benefit from decades of remarkable efforts in the development of numerical methods, algorithms, software, and hardware and witness simulation frontiers that were unimaginable even a few years ago [5]. Large-scale simulations that predict the system’s dynamics may use trillions of computational elements [6] to resolve all spatio-temporal scales, but these often address only idealized systems and their computational cost prevents experimentation and uncertainty quantification.
By contrast, reduced order and coarse-grained models are fast, but limited by the linearization of complex system dynamics while their associated closures, which model the effect of unresolved (not simulated) dynamics on the quantities of interest, are in general based on heuristics and, as a result, domain specific [7]. A closure discovery framework that is independent of the system of interest and can be applied to various domains and tasks is thus highly desirable [8].
To address this challenge we propose Closure-RL, a framework to complement coarse-grained simulations with closures that are discovered by a grid-based Reinforcement Learning (RL) framework. Closure-RL employs a central policy that is based on a FCN and uses locality as an inductive bias. It is able to correct the error of the numerical discretization and improves the overall accuracy of the coarse-grained simulation.
Our approach is inspired by pixelRL [9], a recent work in Reinforcement Learning (RL) for image reconstruction, which minimizes local reconstruction errors. PixelRL employs a per-pixel reward and value network and can therefore be interpreted as a cooperative Multi-Agent Reinforcement Learning (MARL) framework with one agent per pixel. We extend this framework to closure discovery by treating the latter as a reconstruction problem. The numerical scheme introduces corruptions, that the agents are learning to reverse. In contrast to actions based on a set of filters as in pixelRL, we employ a continuous action space that is independent of the PDE to be solved. Our agents learn and act locally on the grid, in a manner that is reminiscent of the numerical discretizations of PDEs based on Taylor series approximations. Hereby, each agent only gets information from its spatial neighborhood.
Although Closure-RL has many characteristics of a cooperative MARL approach, the training costs are similar to a single-agent RL formulation as our proposed formulation employs a central policy and does not require complex interactions between the agents. This allows us to use a large number of agents and no fine-tuning regarding the placement of the agents is required. After training, the framework is capable of accurate predictions in both in- and out-of-distribution test cases. We remark that the actions taken by the agents are highly correlated with the numerical errors and can be viewed as an implicit correction to the effective convolution performed via the coarse-grained discretization of the PDEs [10]. We note that we have chosen a RL-based approach to not require access to a differentiable solver or potentially difficult gradient computations via the adjoint method [8]. Our approach is non-intrusive and can be seamlessly integrated with any numerical discretization scheme.
The main contribution of this work is a grid-based RL algorithm for the discovery of closures for coarse-grained PDEs. The algorithm provides an automated process for the correction of the errors of the associated numerical discretization. We show that the proposed method is able to capture scales beyond the training regime and provides a potent method for solving PDEs with high accuracy and limited resolution.
@@ 2 Related Work
Machine Learning and Partial Differential Equations: In recent years, there has been significant interest in learning the solution of PDEs using Neural Networks. Techniques such as PINNs [11, 12], DeepONet [13], the Fourier Neural Operator [14], NOMAD [15], Clifford Neural Layers [16] and an invertible formulation [17] have shown promising results for both forward and inverse problems. However, there are concerns about their accuracy and related computational cost, especially for low-dimensional problems [18]. These methods aim to substitute numerical discretizations with neural nets, in contrast to our RL framework, which aims to complement them. Moreover, their loss function is required to be differentiable, which is not necessary for the stochastic formulation of the RL reward.
Reinforcement Learning: The present approach is designed to solve various PDEs using a central policy and it is related to similar work for image reconstruction [9]. Its efficient grid-based formulation is in sharp contrast to multi-agent learning formulations that train agents on decoupled subproblems or learn their interactions [19, 20, 21, 22, 23]. The single focus on the local discretization error allows for a general method that is not required to be fine-tuned for the actual application by selecting appropriate global data (such as the energy spectrum) [24] or domain-specific agent placement [25].
@@
Closure Modeling: The development of machine learning methods for discovering closures for under-resolved PDEs has gained attention in recent years. Current approaches are mostly tailored to specific applications such as turbulence modeling [26, 27, 24, 25] and use data such as energy spectra and drag coefficients of the flow in order to train the RL policies. In [28], a more general framework based on diffusion models is used to improve the solutions of Neural Operators for temporal problems using a multi-step refinement process. Their training is based on supervised learning in contrast to the present RL approach which additionally complements existing numerical methods instead of neural network based surrogates [14, 29].
@@
Inductive Bias: The incorporation of prior knowledge regarding the physical processes described by the PDEs, into machine learning algorithms is critical for their training in the Small Data regime and for increasing the accuracy during extrapolative predictions [30]. One way to achieve this is by shaping appropriately the loss function [12, 31, 32, 33, 34], or by incorporating parameterized mappings that are based on known constraints [35, 36, 37]. Our RL framework incorporates locality and is thus consistent with numerical discretizations that rely on local Taylor series based approximations. The incorporation of inductive bias in RL has also been attempted by focusing on the beginning of the training phase [38, 39] in order to shorten the exploration phase.
@@ 3 Methodology
We consider a time-dependent, parametric, non-linear PDE defined on a regular domain $\Omega$. The solution of the PDE depends on its initial conditions (ICs) as well as its PDE-parameters (PDEP) $C\in\mathcal{C}$. The PDE is discretized on a spatiotemporal grid and the resulting discrete set of equations is solved using numerical methods.
In turn, the number of the deployed computational elements and the structure of the PDE determine whether all of its scales have been resolved or whether the discretization amounts to a coarse-grained representation of the PDE. In the first case, the fine grid simulation (FGS) provides the discretized solution $\boldsymbol{\psi}$, whereas in coarse grid simulation (CGS) the resulting approximation is denoted by $\tilde{\boldsymbol{\psi}}$.111 In the following, all variables with the $\;\tilde{}\;$ are referring to the coarse-grid description. The RL policy can improve the accuracy of the solution $\tilde{\boldsymbol{\psi}}$ by introducing an appropriate forcing term in the right-hand side of the CGS. For this purpose, FGS of the PDE are used as training episodes and serve as the ground truth to facilitate a reward signal. The CGS and FGS employed herein are introduced in the next section. The proposed grid-based RL framework is introduced in Section 3.2.
@@ 3.1 Coarse and Fine Grid Simulation
Similarly, we define the time stepping operator of the FGS as $\mathcal{F}:\Psi\times\mathcal{C}\rightarrow\Psi$. In the case of the CGS, $\tilde{C}\in\tilde{\mathcal{C}}$, is an adapted version of $C$, which for instance involves evaluating a PDE-parameter function using the coarse grid.
To simplify the notation, we set $n:={\tilde{n}}=d_{t}n_{f}$ in the following. The update rule of the FGS, i.e. applying $\mathcal{F}$ $d_{T}$-times to advance for the same time increment then the CGS, is referred to as
@@
In Section 4.1 and Section 4.2 we are introducing the FGS and CGS for our illustrative examples. In line with our experiments and to further simplify notation, we are dropping the third spatial dimension in the following presentation.
@@ 3.2 RL Environment
The environment of our RL framework is summarized in Figure 1. We define the state at step $n$ of the RL environment as the tuple $\boldsymbol{S}^{n}:=(\boldsymbol{\psi}^{n},\tilde{\boldsymbol{\psi}^{n}})$. This state is only partially observable as the policy is acting only in the CGS. The observation $\boldsymbol{O}^{n}:=(\tilde{\boldsymbol{\psi}^{n}},\tilde{C}^{n})\in\mathcal{O}$ is defined as the coarse representation of the discretized solution function and the PDEPs. Our goal is to train a policy $\pi$ that makes the dynamics of the CGS to be close to the dynamics of the FGS. To achieve this goal, the action ${\boldsymbol{A}^{n}}\in\mathcal{A}:=\mathbb{R}^{k\times\tilde{N}_{x}\times\tilde{N}_{y}}$ at step $n$ of the environment is a collection of forcing terms for each discretization point of the CGS. In case the policy is later used to complement the CGS simulation the update function in Equation 1 changes to
@@
To encourage learning a policy that represents the non-resolved spatio-temporal scales, the reward is based on the difference between the CGS and FGS at time step $n$. In more detail, we define a local reward $\boldsymbol{R}^{n}\in\mathbb{R}^{\tilde{N}_{x}\times\tilde{N}_{y}}$ inspired by the reward proposed for image reconstruction in [9]:
@@
3.3 Grid-based RL Formulation
The policy $\pi$ predicts a local action $\boldsymbol{A}^{n}_{i,j}\in\mathbb{R}^{k}$ at each discretization point which implies a very high dimensional continuous action space. Hence, formulating the closures with a single agent is very challenging. However, since the rewards are designed to be inherently local, locality can be used as inductive bias and the RL learning framework can be interpreted as a multi-agent problem [9]. One agent is placed at each discretization point of the coarse grid with a corresponding local reward $R_{ij}^{n}$. We remark that this approach augments adaptivity as one can place extra agents at additional, suitably selected, discretization points.
Each agent develops its own optimal policy, which we later defined to be shared, and Equation 5 is replaced by
$\pi_{ij}^{*}=\underset{\pi_{ij}}{\operatorname{argmax}}\ \mathbb{E}_{\pi_{ij}}\left(\sum_{n=0}^{\infty}\gamma^{n}R_{ij}^{n}\right),$ (7)
Here, we used $\mathcal{O}(\tilde{N}_{x}\tilde{N}_{y})$ agents, which for typically used grid sizes of numerical simulations, becomes a large number compared to typical MARL problem settings [22].
We parametrize the local policies using neural networks. However, since training this many individual neural nets can become computationally expensive, we parametrize all the agents together using one fully convolutional network (FCN) [40], which implies that the agents share one policy.
@@ 3.4 Parallelizing Local Policies with a FCN
We define the collection of policies in the FCN as $\Pi^{FCN}:\mathcal{O}\rightarrow\mathcal{A}$. In further discussions, we will refer to $\Pi^{FCN}$ as the global policy. The policy $\pi_{ij}$ of the agent at point $(i,j)$ is subsequently implicitly defined through the global policy as $\pi_{ij}(O_{ij}):=\left[\Pi^{FCN}(\boldsymbol{O})\right]_{:ij}.$ Here, $\boldsymbol{O}_{ij}$ contains only a part of the input contained in $\boldsymbol{O}$  [^2] The exact content of $\boldsymbol{O}_{ij}$ is depending on the receptive field of the FCN. For consistency, we will refer to $\boldsymbol{O}_{ij}$ as a local observation and $\boldsymbol{O}$ as the global observation. We note that the policies $\pi_{ij}(\boldsymbol{O}_{ij})$ map the observation to a probability distribution at each discretization point (see Appendix A for details including the neural network architecure employed). Similar to the global policy, the global value function is parametrized using a FCN as well. It maps the global observation to an expected return $H\in\mathcal{H}:=R^{\tilde{N}_{x}\times\tilde{N}_{y}}$ at each discretization point $\mathcal{V}^{FCN}:\mathcal{O}\rightarrow\mathcal{H}.$ Similarly to the local policies, we define the local value function related to the agent at point $(i,j)$ as $v_{ij}(O_{ij}):=[\mathcal{V}^{FCN}(\boldsymbol{O})]_{ij}.$
3.5 Policy Optimization
In order to solve the optimization problem in Equation 7, we employ a modified version of the PPO algorithm [41].
@@ 3.5 Policy Optimization
$\mathbb{E}_{\boldsymbol{S},\boldsymbol{A}\sim\Pi^{FCN}}\left[\frac{1}{\tilde{N}_{x}\cdot\tilde{N}_{y}}\sum_{i,j=1}^{\tilde{N}_{x},\tilde{N}_{y}}L_{\pi_{ij}}(\boldsymbol{O}_{ij},\boldsymbol{A}_{ij})\right]$ (8)
with the local version of the PPO objective $L_{\pi_{ij}}(\boldsymbol{O}_{ij},\boldsymbol{A}_{ij})$. This corresponds to the local objective of the policy of the agent at point $(i,j)$ [41].
@@
We have provided an overview of the modified PPO algorithm in Appendix B together with further details regarding the local version of the PPO objective $L_{\pi_{ij}}(\boldsymbol{O}_{ij},\boldsymbol{A}_{ij})$. Our implementation of the adapted PPO algorithm is based on the single agent PPO algorithm of the Tianshou framework [42].